# SADA: Semantic adversarial unsupervised domain adaptation for Temporal Action Localization

### 20 Dec 2023 v1

@@ Abstract
Temporal Action Localization (TAL) is a complex task that poses relevant challenges, particularly when attempting to generalize on new – unseen – domains in real-world applications. These scenarios, despite realistic, are often neglected in the literature, exposing these solutions to important performance degradation. In this work, we tackle this issue by introducing, for the first time, an approach for Unsupervised Domain Adaptation (UDA) in sparse TAL, which we refer to as Semantic Adversarial unsupervised Domain Adaptation (SADA). Our contribution is threefold: (1) we pioneer the development of a domain adaptation model that operates on realistic sparse action detection benchmarks; (2) we tackle the limitations of global-distribution alignment techniques by introducing a novel adversarial loss that is sensitive to local class distributions, ensuring finer-grained adaptation; and (3) we present a novel experimental setup, based on EpicKitchens100, that evaluates multiple types of domain shifts in a comprehensive manner. Our experimental results indicate that SADA improves the adaptation across domains when compared to fully supervised state-of-the-art and alternative UDA methods, attaining a relative performance boost of up to .
1 Introduction
Recent advances in the field of video understanding have played a critical role in the surge of novel video-based applications – e.g., video indexing, summarization, recommendation, or surveillance. A critical task of this field is Temporal Action Localization (TAL), which involves identifying actions in a video consisting of both their time intervals and action categories. This is particularly difficult given the inherent variabilities of videos. These can be presented, among others, in the form of appearance variability – e.g. induced by differences in the kitchen – or as acquisition variability – e.g. different acquisition devices, lighting conditions, or acquisition viewpoints. All in all, these prompt a certain degree of confusion between similar actions to be discriminated.
@@ 1 Introduction
Unsupervised Domain Adaptation (UDA) has recently become a hot topic given its potential to leverage unlabelled data to mitigate this domain-induced degradation [17, 41, 21]. Unfortunately, despite its considerable success in image-based tasks [11], the application of UDA to video understanding remains underexplored. In fact, to the best of our knowledge, no prior work addresses UDA for TAL setups. The closest proposal is SSTDA [8], which approaches the problem of action segmentation. This focuses on making per-frame (namely dense) action predictions, tackling datasets where no concurrent actions take place [15, 29]. Hence, not being able to deal with sparse detections nor simultaneous intersection of labels.
@@
Consequently, in this paper, we propose the first Unsupervised Domain Adaptation method for sparse detection on TAL, which we name Semantic Adversarial unsupervised Domain Adaptation or SADA for short. Our goal is to minimize the discrepancy between feature representations of a labeled source domain and an unlabelled target domain. These features are extracted with a multi-resolution architecture [49] that we couple with a novel adversarial loss that improves the limitations of existing UDA works. Concretely, existing works normally align domain distributions globally [17]. We propose instead to use pseudo-labeling [30] to assign an action or background (no action) class to each feature representation. With this, our loss factorizes the global adaptation strategy into independent per-class and background alignments – i.e., aligning each action’s distribution across both domains. This results in a more sensitive alignment strategy compared to a global distribution approach, less prone to feature misalignment and, as we will show, better performing.
@@
Assessing the effectiveness of unsupervised domain adaptation methods for video understanding is a challenging, still unresolved task. Existing proposals on action segmentation [8] follow a subject-based strategy where they aim to adapt a model to new unseen subjects. Here we refer to subject as a person appearing in a video. Nevertheless, little data is normally available from a single subject, which inevitably requires grouping several of them for training. This allows the model to generalize over the subject variability under study, making it unsuitable to test for domain adaptation. We overcome these limitations by proposing a novel comprehensive suite of 6 new benchmarks for sparse TAL based on EpicKitchens100 [14] dataset. These study appearance and acquisition domain shifts, yielding a fine-grained and reliable evaluation of the model’s ability to operate on new unseen domains. This comprehensive benchmark allows us to demonstrate that our approach mitigates the performance degradation, improving by a large margin the existing fully supervised (namely source-only) and unsupervised domain adaptation-based proposals.
In short, our main contributions are three-fold:
@@
We present a new set of benchmarks to test sparse detection scenarios when facing 6 different domain shifts, achieving state-of-the-art results in all of them.
@@ 2 Related work
Unsupervised Domain Adaptation. Domain Adaptation techniques emerge as an effective solution to bridge the gap between data collected from a source and a target distribution, respectively. A large suite of approaches has been proposed to perform this alignment between the labeled and the unlabeled domains – e.g., discrepancy minimization [28, 60], entropy minimization [19, 59], contrastive learning [26] or domain normalization [56]. The most popular approach is arguably the use of adversarial training methods [17, 41, 22, 54, 38, 21], which incorporate a domain classifier trained to discern if the samples come from the source or the target domain, in a min-max fashion. This results in the computation of domain-invariant embeddings [12]. Despite convenient, the simplicity of these methods might degrade the quality of the alignment [30], as they potentially align embeddings of source and target domain that represent different semantic information – e.g., different class labels. Few works have been proposed to do this alignment in a more sensitive way [30]. We highlight the importance of [58], proposing to minimize instead the distance of per-class centroids, computed as the mean of the set of predicted feature embeddings of a given class. Its effectiveness, however, relies on the assumption that the data is distributed somewhat homogeneously around the center, as otherwise, the centroids are not necessarily meaningful. In our work, we couple the advantages of both adversarial domain adaptation and semantic alignment and propose for the first time a pure adversarial semantic loss that yields domain invariant representations in a semantically meaningful way, without making explicit assumptions of the distributions (see Fig. 1).
Unsupervised Domain Adaptation for TAL: Despite the considerable success of UDA methods, their applicability has been mostly restricted to image-based scenarios [11] such as image classification [20, 17, 37, 41, 48], object detection [9, 42] or semantic segmentation [47, 23]. Much less attention has been dedicated to video-based applications such as action recognition [7, 25, 43] or spatio-temporal action segmentation [1, 40]. To the best of our knowledge, at the time of this writing, there is no direct comparison with our work focusing on sparse Temporal Action Localization. The closest work is SSTDA [8] that applies unsupervised domain adaptation for Action Segmentation. SSTDA proposes the use of two global-distribution-based auxiliary tasks to jointly align cross-domain feature spaces. Unlike our proposal, their work falls into the category of anchor-free, making per-frame action predictions. This restricts its applicability to action segmentation scenarios, where current datasets [15, 29] are designed to deal with frame-based single-action classification. In our work, we overcome this limitation by leveraging an anchor-based architecture that enables a natural adaptation to more realistic multi-label scenarios.
@@ 3.3 Backbone and SGP pyramid
Our model first takes two input videos and of both domains and . For simplicity, both videos and have length , which we enforce using padding. The method then processes the two raw videos applying a frozen pre-trained backbone – e.g., I3D [5], Slowfast [16]. This permits to extract, in an effective way, temporal cues of the video into a set of refined video features. These embeddings are then fed to an SGP feature pyramid [49] which combines the use of SGP blocks and the progressive downsampling of the temporal length by a ratio of . This outputs a set of multi-resolution anchor embeddings and , for the two domains, respectively. Here denotes the set of predefined resolution levels and the anchor embeddings of level of a given domain. Concretely, this permits to obtain embeddings for a set of uniformly sampled anchors at each of the resolution levels. The use of a multi-resolution model is favorable to naturally adapt to different action lengths.
@@ 3.4 Classification and localization head
To learn the underlying TAL task, we train in a fully supervised manner a classification and localization module with the labeled source domain . Due to the anchor-based nature of our model, we first require a matching strategy between the set of candidate anchors to the actual GT segments. For this, we follow a center sampling strategy [53, 63, 62, 49]. In other words, for a given level , we define an anchor as positive if the time instant that it represents is near the center of an action. All the rest are marked as negative. We define and as the begins, ends and action classes of their matching GT segments. We identify negative anchors by assigning them to action class 0.
@@
Here and are tunable parameters that control the importance of each of the two losses, respectively.
@@ 3.5 Our proposal: Semantic adversarial multi-resolution alignment
One of the main contributions of this paper is the design of a novel adversarial-based loss that we name SADA loss, which attempts to overcome the limitations of the extensively used global adversarial loss [17]. Traditional adversarial domain adaptation relies on the idea of designing a domain classifier that learns to identify the domain that each of the embeddings belongs to. The rest of the model learns concurrently the opposite objective, which, as shown by [17], results in the learning of the domain invariant representations. While this approach has been shown to be effective in other fields – e.g., image classification, object detection – we find that its performance in a more challenging setup like TAL still presents important challenges. One of the main issues, as argued by [30] is that this loss often suffers from feature misalignment which greatly declines the model’s performance. This refers to the cases where these methods align embeddings of non-matching class labels – e.g., aligning embeddings of domain of an action with embeddings of domain of a class .
3.5.1 Local adversarial alignment
To fix this feature misalignment in realistic scenarios like TAL, we propose an alternative adversarial loss formulation that provides a finer-grained alignment. This loss first attempts to perform a local class-aware alignment. This is, for every given resolution level , we group the anchor embeddings of and corresponding to an action label . This is straightforward for the source domain as we have the GT information. For the target domain, we use a hard-pseudo labeling strategy [30] that classifies a given embedding as class if this is the highest-confidence score of the predicted class distribution, and this is above a threshold . Formally we define the hard pseudo-label of a given anchor as:
@@ 3.5.1 Local adversarial alignment
where and are the labels of each of the domains. Following [17], we introduce a Reverse Gradient Layer (GRL) before the discriminator to invert the gradients sign which creates the min-max game where the feature extractor learns to confuse the discriminator. We condition the discriminator to class using a learnable class embedding that we replicate for every selected anchor into an embedding . See the Supp. for an ablation that justifies this design decision.
3.5.2 Local and global alignment: SADA
Eq. 7 aligns solely the anchor embeddings that are classified as one of the classes, but what happens with the background embeddings that fall below the threshold ? In this case, the loss completely ignores their influence, yielding a suboptimal partial alignment – as shown in Sec. 4.
@@ 4 Datasets and experiments
In this section, we present a novel comprehensive benchmark to evaluate the task of UDA for TAL. Our setup, for the first time, goes beyond action segmentation and evaluates the adaptation to different domain shifts in more realistic scenarios with sparse multi-label annotations. We then showcase the effectiveness and competitiveness of our model over the state-of-the-art methods together with several relevant ablations.
4.1 Experimental setup
Evaluating domain adaptation-based methods in the context of video understanding is a challenging issue that requires the definition of a reasonable domain gap and identifying a sufficiently large set of intersecting action classes. Dividing existing datasets into different domains that comply with these conditions often restricts the amount of data to learn and adapt. SSTDA [8] approaches this problem on GTEA [15] and Breakfast [29] by defining a subject-based partitioning where they aim to adapt the model to new unseen subjects. However, as little data is available from a single subject in those datasets, they group several users for training. This allows the model to generalize over the subject variability under study, making it unsuitable to test for domain adaptation. Closely related to our work, TranSVAE [57] defines a source-target for video classification based on EpicKitchens100 [14]. They split the data into three different kitchens as different domains and perform cross-kitchen evaluation, limiting even further the amount of data in each domain.
A new benchmark for UDA on sparse TAL. To overcome these limitations, we propose a set of 6 different benchmarks for sparse TAL based on the EpicKitchen100 [14]. EpicKitchen100 presents an ideal base for our tasks as it has become a gold standard to evaluate complex sparse detection scenarios on long egocentric videos (up to 45 minutes). We identify two domain gaps in this dataset: an appearance domain shift based on the different colors of the kitchen counters; and an acquisition domain shift that results in the differences of lighting and camera conditions when extending EpicKitchens55 [14] into its new version EpicKitchens100 [14]. This permits to define a rich set of benchmarks that provide a more fine-grained and comprehensive evaluation. Importantly, this strategy is also suitable for single-source settings which do not allow an easy generalization over the shift under study. For example, if a model is trained with dark-counter kitchens only, it cannot easily generalize to white-counter kitchens.
@@ 4.1 Experimental setup
One final important consideration is that EpicKitchens100 [14] exhibits a very strong long-tail label distribution, where the vast majority of its 97 actions represent only a marginal percentage of the overall data. Adapting domain adaptation methods to these long tail distributions falls beyond the scope of this paper and most of the existing literature. For this reason, we limit our evaluation to the 10 majority classes representing of the original dataset. This still results in a very complex task proof of which is that the best-performing models to date attain less than a 30% mAP [62, 49].

### 05 Apr 2024 v2

@@ Abstract
Temporal Action Localization (TAL) is a complex task that poses relevant challenges, particularly when attempting to generalize on new – unseen – domains in real-world applications. These scenarios, despite realistic, are often neglected in the literature, exposing these solutions to important performance degradation. In this work, we tackle this issue by introducing, for the first time, an approach for Unsupervised Domain Adaptation (UDA) in sparse TAL, which we refer to as Semantic Adversarial unsupervised Domain Adaptation (SADA). Our contributions are threefold: (1) we pioneer the development of a domain adaptation model that operates on realistic sparse action detection benchmarks; (2) we tackle the limitations of global-distribution alignment techniques by introducing a novel adversarial loss that is sensitive to local class distributions, ensuring finer-grained adaptation; and (3) we present a novel set of benchmarks based on EpicKitchens100 and CharadesEgo, that evaluate multiple domain shifts in a comprehensive manner. Our experiments indicate that SADA improves the adaptation across domains when compared to fully supervised state-of-the-art and alternative UDA methods, attaining a performance boost of up to mAP.
Index Terms:
Artificial Intelligence, Computer Vision, Video Understanding, Domain Adaptation1 Introduction
Recent advances in the field of video understanding have played a critical role in the surge of novel video-based applications –e.g., video indexing, summarization, recommendation, or surveillance. A critical task of this field is Temporal Action Localization (TAL), which involves identifying actions in a video consisting of both their time intervals and action categories. This is particularly difficult given the inherent variabilities of videos. These can be presented, among others, in the form of appearance variability –e.g., different kitchens and/or lighting conditions–, acquisition variability –e.g., different recording devices– or viewpoint variability –e.g., first- or third-person. All in all, these prompt a certain degree of confusion between similar actions to be discriminated.
@@
Unsupervised Domain Adaptation (UDA) has recently become a hot topic given its potential to leverage unlabelled data to mitigate this domain-induced degradation [3, 4, 5]. Unfortunately, despite its considerable success in image-based tasks [6], the application of UDA to video understanding remains underexplored. In fact, to the best of our knowledge, no prior work addresses UDA for TAL setups. The closest proposal is SSTDA [7], which approaches the problem of action segmentation. This focuses on making per-frame action predictions (i.e., dense TAL), tackling datasets where no concurrent actions take place [8, 9]. This approach inevitably requires additional smoothing techniques to preserve temporal coherence. This motivates the focus of this paper on sparse TAL –i.e., segment-level predictions– avoiding the need for additional losses, while intrinsically adapting to multi-label scenarios.
@@
Consequently, in this paper, we propose the first Unsupervised Domain Adaptation method for sparse multi-label detection on TAL, which we name Semantic Adversarial unsupervised Domain Adaptation or SADA for short. Our goal is to minimize the discrepancy between feature representations of a labeled source domain and an unlabelled target domain. These features are extracted with a multi-resolution architecture [10] that we couple with a novel adversarial loss that improves the limitations of existing UDA works. Concretely, existing works normally align domain distributions globally [3]. We propose instead to use pseudo-labeling [11] to assign an action or background (no action) class to each feature representation. With this, our loss factorizes the global adaptation strategy into independent per-class and background alignments – i.e., aligning each action’s distribution across both domains. This results in a more sensitive alignment strategy compared to a global distribution approach, less prone to feature misalignment and, as we will show, better performing.
@@
Assessing the effectiveness of UDA methods for video understanding is a challenging, still unresolved task. Existing proposals on action segmentation [7] follow a subject-based strategy where they aim to adapt a model to new unseen subjects. Here we refer to subject as a person appearing in a video. Nevertheless, little data is normally available from a single subject, which inevitably requires grouping several of them for training. This allows the model to generalize over the subject variability under study, making it unsuitable for domain adaptation. To address these limitations, we draw inspiration from the work of [12] on action recognition and investigate the impact of viewpoint domain shifts on sparse TAL in CharadesEgo [13]. However, we contend that a more comprehensive evaluation necessitates setups with more controllable shifts. For this, we also propose a suite of 6 new setups based on EpicKitchens100 [14] which study the effect appearance and acquisition domain shifts. These comprehensive benchmarks demonstrate that SADA mitigates the performance degradation, improving by a large margin the existing fully supervised (namely source-only) and UDA-based proposals. In short, our main contributions are:
@@
We present new benchmarks to test sparse detection scenarios when facing 7 different domain shifts, improving the state-of-the-art in all of them.
@@ 2 Related work
Unsupervised Domain Adaptation. Domain Adaptation techniques emerge as an effective solution to bridge the gap between data collected from a source and a target distribution, respectively. A large suite of approaches has been proposed to perform this alignment between labeled and unlabeled domains – e.g., discrepancy minimization [34, 35], entropy minimization [36, 37] or contrastive learning [38]. The most popular approach is arguably the use of adversarial training methods [3, 4, 39, 40, 41, 42, 5], which incorporate a domain classifier trained to discern if samples come from the source or the target domain, in a min-max fashion. This results in the computation of domain-invariant embeddings [43]. Despite convenient, the simplicity of these methods might degrade the quality of the alignment [11], as they potentially align embeddings of source and target domain that represent different semantic information – e.g., different class labels. Few works have been proposed to do this alignment in a more sensitive way [11]. We highlight the importance of [2], proposing to minimize instead the distance of per-class centroids, computed as the mean of the set of predicted feature embeddings of a given class. Its effectiveness, however, relies on the assumption that the data is distributed somewhat homogeneously around the center, as otherwise, the centroids are not necessarily meaningful. In our work, we couple the advantages of both adversarial domain adaptation and semantic alignment and propose for the first time a pure adversarial semantic loss that yields domain invariant representations in a semantically meaningful way, without making explicit assumptions of the distributions (see Fig. 1).
Unsupervised Domain Adaptation for TAL: Despite the considerable success of UDA methods, their applicability has been mostly restricted to image-based scenarios [6] such as image classification [44, 3, 45, 4, 46], object detection [47, 48] or semantic segmentation [49, 50]. Much less attention has been dedicated to video-based applications such as action recognition [51, 52, 53] or spatio-temporal action segmentation [54, 55]. To the best of our knowledge, at the time of this writing, there is no direct comparison with our work focusing on UDA for sparse Temporal Action Localization. The closest work is SSTDA [7] that applies unsupervised domain adaptation for Action Segmentation. SSTDA proposes the use of two global-distribution-based auxiliary tasks to jointly align cross-domain feature spaces. Unlike our proposal, their work falls into the category of anchor-free, making per-frame action predictions. This restricts its applicability to action segmentation scenarios, where current datasets [8, 9] are designed to deal with frame-based single-action classification. In our work, we overcome this limitation by leveraging an anchor-based architecture that enables a natural adaptation to more realistic multi-label scenarios.
@@ 3.3 Backbone and SGP pyramid
Our model first takes two input videos and of both domains and . For simplicity, both videos and have length , which we enforce using padding. The method then processes the two videos applying a frozen pre-trained backbone – e.g., I3D [56], Slowfast [57]. This permits to extract, in an effective way, temporal cues of the video into a set of refined video features. These embeddings are then fed to an SGP feature pyramid [10] which combines the use of SGP blocks and the progressive downsampling of the temporal length by a ratio of . This outputs a set of multi-resolution anchor embeddings and , for the two domains, respectively. Here denotes the set of predefined resolution levels and the anchor embeddings of level of a given domain. Concretely, this permits to obtain embeddings for a set of uniformly sampled anchors at each of the resolution levels. The use of a multi-resolution model is favorable to naturally adapt to different action lengths.
@@ 3.4 Classification and localization head
To learn the underlying TAL task, we train in a fully supervised manner a classification and localization module with the labeled source domain . Due to the anchor-based nature of our model, we first require a matching strategy between the set of candidate anchors to the actual GT segments. For this, we follow a center sampling strategy [58, 59, 21, 10]. In other words, for a given level , we define an anchor as positive if the time instant that it represents is near the center of an action. All the rest are marked as negative. We define and as the begins, ends, and action classes of their matching GT segments. We identify negative anchors by assigning them to action class 0.
@@
Here and are two tunable hyperparameters.
@@ 3.5 Our proposal: Semantic adversarial multi-resolution alignment
One of the main contributions of this paper is the design of a novel adversarial-based loss that we name SADA loss, which attempts to overcome the limitations of the extensively used global adversarial loss [3]. Traditional adversarial domain adaptation relies on the idea of designing a domain classifier that learns to identify the domain that each of the embeddings belongs to. The rest of the model learns concurrently the opposite objective, which, as shown by [3], results in the learning of the domain invariant representations. While this approach has been shown to be effective in other fields – e.g., image classification, and object detection – we find that its performance in a more challenging setup like TAL still presents important challenges. One of the main issues, as argued by [11] is that this loss often suffers from feature misalignment which greatly declines the model’s performance. This refers to the cases where these methods align embeddings of non-matching class labels – e.g., aligning embeddings of domain of an action with embeddings of domain of a class .
Local adversarial alignment: To fix this feature misalignment in realistic scenarios like TAL, we propose an alternative adversarial loss formulation that provides a finer-grained alignment. This loss first attempts to perform a local class-aware alignment. This is, for every given resolution level , we group the anchor embeddings of and corresponding to an action label . This is straightforward for the source domain as we have the GT information. For the target domain, we use a hard-pseudo labeling strategy [11] that classifies a given embedding as class if this is the highest-confidence score of the predicted class distribution, and this is above a threshold . Formally we define the hard pseudo-label of a given anchor as:
@@
where and are the labels of each of the domains. Following [3], we introduce a Reverse Gradient Layer (GRL) before the discriminator to invert the sign of the gradient which creates the min-max game where the feature extractor learns to confuse the discriminator. We condition the discriminator to class using a learnable class embedding that we replicate for every selected anchor into an embedding .
Local and global alignment (SADA): Eq. 7 aligns solely the anchor embeddings that are classified as one of the classes, but what happens with the background embeddings that fall below the threshold ? In this case, the loss ignores their influence, yielding a suboptimal partial alignment (see Table IV).
@@ 4 Datasets and experiments
In this section, we present a novel comprehensive benchmark to evaluate the task of UDA for TAL. Our setup, for the first time, goes beyond action segmentation and evaluates the adaptation to different domain shifts in more realistic scenarios with sparse multi-label annotations. We then showcase the effectiveness of our model over the state-of-the-art methods together with several relevant ablations.
4.1 Benchmarks for UDA on sparse TAL
Evaluating domain adaptation-based methods in the context of video understanding is a challenging issue that requires the definition of a reasonable domain gap and identifying a sufficiently large set of intersecting action classes. Dividing existing datasets into different domains that comply with these conditions often restricts the amount of data to learn and adapt. SSTDA [7] approaches this problem on GTEA [8] and Breakfast [9] by defining a subject-based partitioning where they aim to adapt the model to new unseen subjects. However, as little data is available from a single subject in those datasets, they group several users for training. This allows the model to generalize over the subject variability under study, making it unsuitable to test for domain adaptation. Closely related to our work, [61, munro2020multi] propose several UDA scenarios for video classification based on EpicKitchens100 [14]. These define 3 domains based on the data of 3 different kitchens, thus performing cross-kitchen evaluation. This limits even further the amount of data in each domain –i.e., between 15 to 29 videos per domain.
EpicKitchens100: To overcome these limitations, we first propose a new set of 6 different scenarios (S1, …, S6) for sparse TAL based on the EpicKitchen100 [14] (see Fig. 3). EpicKitchen100 presents an ideal base for our tasks as it has become a gold standard to evaluate complex sparse detection scenarios on long egocentric videos (up to 45 minutes). We identify two domain gaps in this dataset: an appearance domain shift based on the different colors of the kitchen counters; and an acquisition domain shift that results in the differences of lighting and camera conditions when extending EpicKitchens55 [14] into its new version EpicKitchens100 [14]. This permits to define a rich set of benchmarks that provide a more fine-grained and comprehensive evaluation. Importantly, this strategy is also suitable for single-source settings which do not allow an easy generalization over the shift under study. For example, if a model is trained with dark-counter kitchens only, it cannot easily generalize to white-counter kitchens.
@@ 4.1 Benchmarks for UDA on sparse TAL
One important consideration is that EpicKitchens100 exhibits a very strong long-tail label distribution, where the vast majority of its 97 actions represent only a marginal percentage of the overall data (see Supp.). Adapting DA methods to these long tail distributions falls beyond the scope of this paper and most of the existing literature. For this reason, we follow a similar approach to that proposed by [munro2020multi] and limit our evaluation to the 10 majority classes representing of the original dataset. This still results in a very complex task proof of which is that the best-performing models to date attain less than a 30% mAP [21, 10].
CharadesEgo: One important aspect of the 6 previous setups is that all the domains share an egocentric viewpoint. For this reason, we evaluate the model degradation under extreme domain shifts caused by major changes in the perspective. For this, we use CharadesEgo [13] which extends the original third-person videos from Charades [62] by matching them with their corresponding egocentric videos. Following [12], we define the Source domain as the third-person videos, and the Target domain as the egocentric ones. Given the extreme shift that these present (see Fig. 4), we limit the task to predicting the verb of each of the action segments, and similarly to the previous case, we keep the 10 majority verbs only. This mitigates the effect of long-tail action distributions in our evaluation which falls beyond the scope of this paper. See the Supp. for more details.